# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v1

@@ 5.6.4 Graph Neural Network-Based Methods
Representative architectures of GNNs include graph convolutional network (GCN) (Kipf and Welling, 2017), graph attention network GAT) (Veličković et al., 2017), graph isomorphism network (GIN) (Xu et al., 2019a), and graph transformer (GT) (Yun et al., 2019). The hidden layers of a GCN can be formulated as
@@
Several studies have explored using Graph Neural Networks (GNNs) for missing data imputation (Zhang and Chen, 2020; You et al., 2020; Spinelli et al., 2020; Telyatnikov and Scardapane, 2023). A key challenge, however, is that GNNs typically require a pre-defined similarity graph across the entire dataset, which is difficult to construct from incomplete data. Zhang and Chen (2020) proposed a GNN-based matrix completion method for collaborative filtering. You et al. (2020) proposed GRAPE, a graph-based framework for feature imputation as well as label prediction. GRAPE tackles the missing data problem using a graph representation, where the observations and features are viewed as two types of nodes in a bipartite graph, and the observed feature values as edges. Telyatnikov and Scardapane (2023) proposed EGG-GAE, a GCN-based method for tabular data that adaptively constructs the adjacency matrix from a neural network’s representations of the incomplete data. In a different domain, Li et al. (2024b) combined a GNN with a diffusion model to impute missing values in spatial transcriptomics data. More methods of GNN-based imputation can be found in (Huang et al., 2022; Wu et al., 2023; Eskandari et al., 2024; Zhang et al., 2025e).
@@ 5.7 Large Language Model for Missing Data Imputation
Large Language Models (LLMs) (Achiam et al., 2023; Naveed et al., 2025) are billion-parameter neural networks that learn universal linguistic patterns by next-token prediction on web-scale text. Their recent ascendancy has redefined state-of-the-art across conversational AI, code synthesis, and scientific discovery. It is also promising to exploit LLMs in missing data imputation tasks. Indeed, the pretrained language models are already effective enough in predicting the missing words in pure text owing to the training strategy of next-token prediction, of which a typical application is code completion (Wei et al., 2023; Izadi et al., 2024; Husein et al., 2025). For non-text data such as tabular data, a rudimentary approach is to feed the LLM an incomplete table as raw text and prompts it to hallucinate the blanks; while clever prompting can yield plausible guesses, this tactic barely scratches the surface of LLMs’ expressive power.
@@
Mei et al. (2021) proposed a semantic-aware imputation framework, called IPM, that leverages pre-trained language models to fill missing categorical data. IPM has two variants: IPM-Multi, which treats imputation as a multiclass classification task, and IPM-Binary, which models it as a binary candidate selection problem to mitigate over-fitting in low-redundancy datasets. By utilizing fine-tuned models like RoBERTa and neighbor-based negative sampling, IPM captures fine-grained token semantics, outperforming traditional and generative baselines across multiple real-world datasets. Zhang et al. (2023) explored the use of LLMs like GPT-4 for data preprocessing tasks such as missing data imputation. The paper introduced a framework that combines prompt engineering techniques—such as zero-shot, few-shot, and batch prompting—with contextualization and feature selection to guide LLMs in imputing missing values. Wang et al. (2024c) introduces TREB, a BERT-based framework for imputing missing values in tabular data. By transforming numerical rows into tokenized text and fine-tuning RoBERTa with a custom 4-digit tokenizer, TREB captures feature interdependencies without relying on categorical semantics. In (Huang and Hao, 2024), the authors proposed a missing data imputation framework called PRPMI. PRPMI enhances pre-trained language models by incorporating trainable prompt embeddings and retrieval-augmented knowledge from external data sources. It serializes tabular data into natural language format, fine-tunes the model using both original and retrieved contextual information, and treats imputation as a multiclass classification task. Ding et al. (2024) proposed to fine-tune a pre-trained GPT-2 model using complete data and prompting it to predict missing values in recommender systems. The method generates semantically meaningful imputations and outperforms a few traditional imputation methods on some benchmarks. Hayat and Hasan (2024) presented a method called CRILM for context-aware missing data imputation that leverages pre-trained language models to generate meaningful, contextually relevant descriptors for missing data instead of relying on traditional numerical estimation. By transforming tabular data into natural language and fine-tuning lighter LLMs on these enriched datasets, CRILM effectively handles various missingness mechanisms (MCAR, MAR, and MNAR), and provides promising imputation performance. Fang et al. (2025b) proposed SPLLM, a spatiotemporal pretrained LLM designed for forecasting with missing values. SPLLM integrates a spatiotemporal fusion GCN module to extract spatial and temporal correlations, employs an FFN fine-tuning strategy to adapt pretrained LLMs to incomplete data, and uses a final fusion layer to combine learned and pretrained knowledge.
@@ 5.8 Other imputation methods
Some imputation methods fall outside the categories discussed in previous sections. For instance, Muzellec et al. (2020) proposed a method based on optimal transport (Villani et al., 2008). Their core idea is that two randomly drawn batches from the same dataset should share the same distribution; this principle forms a loss function for imputation by matching distributions. Building on this, Zhao et al. (2023) proposed imputing missing values by transforming two sample batches into a latent space using deep invertible functions and then matching their distributions. Traditional missing data imputation methods are simple and flexible but require precise model specification, while deep generative approaches are efficient but harder to optimize and rely on stronger assumptions. Motivated by this, Jarrett et al. (2022) introduced HyperImpute, a framework that combines their strengths by adaptively configuring column-wise models. The proposed method is implemented with ready-to-use tools and validated through extensive experiments, demonstrating superior accuracy over benchmarks and reinforcing the effectiveness of iterative imputation.
@@ 6.1 Tensor Completion
Tensors in this paper are referred to as multidimensional arrays (Kolda and Bader, 2009) that, as shown by Figure 9, generalize scalars, vectors, and matrices to higher dimensions. We denote $\bm{\mathcal{T}}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ an order-$d$ or $d$-way real tensor. Higher-order ($d\geq 3$) tensors widely exist in science and engineering. Examples of 3rd-order tensors include RGB images, and hyperspectral images. Examples of 4th-order tensors include a batch of RGB images and videos. Examples of 5th-order tensors include 3D MRI over time and subjects.
@@
Tensor completion aims to recover the missing entries of a tensor satisfying some properties. Let $\Omega$ be the set of locations of observed entries of $\mathcal{T}$. Tensor completion usually solves the following problem
$\mathop{\text{minimize}}_{\bm{\mathcal{X}}}\dfrac{1}{2}\|\mathcal{P}_{\Omega}\left(\bm{\mathcal{T}}-\bm{\mathcal{X}}\right)\|_{F}^{2}+\lambda\cdot R(\bm{\mathcal{X}}),$ (28)
where $R$ denotes a regularizer ensuring a certain property on $\bm{\mathcal{X}}$. Similar to matrix completion, the most common property used in tensor completion is the low-rankness. Low-rank tensor completion (LRTC), as a higher-order extension of LRMC, has been extensively studied in the past decades. A naive approach to LRTC is to unfold a tensor to a matrix first and then perform LRMC or other missing data imputation methods (Liu et al., 2012). However, unfolding will destroy the higher-order structure of tensors and lead to lower imputation performance. Yuan and Zhang (2016) pointed out that the unfolding-based tensor completion algorithms require much more observed entries for exact recovery than direct tensor completion without matricization.
LRTC methods can be organized into different categories according to the types of factorization models. Representative tensor factorization models include CP (CANDECOMP/PARAFAC) decomposition (Carroll and Chang, 1970), Tucker decomposition Tucker (1966), t-SVD decomposition (Kilmer and Martin, 2011), tensor train decomposition (Oseledets, 2011), and tensor ring decomposition (Zhao et al., 2016). In this paper, we take the CP decomposition of a tensor $\bm{\mathcal{T}}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ as an example:
$\bm{\mathcal{T}}\approx\sum_{i=1}^{r}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\circ\cdots\circ\mathbf{a}_{i}^{(d)}$ (29)
@@
$\textup{rank}(\bm{\mathcal{T}})=\min\left\{r\in\mathbb{N}:\ \mathbf{\mathcal{T}}=\sum_{i=1}^{r}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\cdots\circ\mathbf{a}_{i}^{(d)}\right\}.$
@@
CP decomposition based LRTC methods (Acar et al., 2011; Jain and Oh, 2014; Zhao et al., 2015; Liu and Moitra, 2020; Fan et al., 2023) often solve
$\mathop{\text{minimize}}_{\{\mathbf{a}_{i}^{(j)}\}}\dfrac{1}{2}\left\|\mathcal{P}_{\Omega}\left(\bm{\mathcal{T}}-\sum_{i=1}^{r}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\circ\cdots\circ\mathbf{a}_{i}^{(d)}\right)\right\|_{F}^{2}+\lambda\cdot R\left(\{\mathbf{a}_{i}^{(j)}\}\right)$ (30)
The nuclear norm (Friedland and Lim, 2018) of a tensor $\bm{\mathcal{X}}$, though still NP-hard to compute, is defined as
$\|\bm{\mathcal{X}}\|_{\ast}=\inf\left\{\sum_{i=1}^{r}|s_{i}|:\ \bm{\mathcal{X}}=\sum_{i=1}^{r}s_{i}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\cdots\circ\mathbf{a}_{i}^{(d)},\|\mathbf{a}_{i}^{(j)}\|=1,\ r\in\mathbb{N}\right\}.$ (31)
@@
$\mathop{\text{minimize}}_{\{s_{i},\mathbf{a}_{i}^{(j)}\}}\dfrac{1}{2}\left\|\mathcal{P}_{\Omega}\left(\bm{\mathcal{T}}-\sum_{i=1}^{r}s_{i}\cdot\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\circ\cdots\circ\mathbf{a}_{i}^{(d)}\right)\right\|_{F}^{2}+\lambda\cdot\|\mathbf{s}\|_{1}$ (32)
@@
where $\|\mathbf{a}_{i}^{(j)}\|=1$, $i=1,\ldots,r$, $j=1,\ldots,d$. Note that the $\ell_{1}$ norm in (32) can be replaced by $\ell_{p}$ quasi-norms, leading to tensor Schatten-$p$ quasi-norm regularized LRTC (Fan et al., 2023). In (Bazerque et al., 2013; Cheng et al., 2016; Yang et al., 2016; Lacroix et al., 2018; Fan et al., 2023), $s_{i}$ are absorbed into the factor vectors $\mathbf{a}_{i}^{(j)}$, which makes the optimization more efficient. In particular, (Fan et al., 2023) presented a class of tensor rank regularizers based on the Euclidean norms of the CP component vectors of a tensor and show that these regularizers are monotonic transformations of tensor Schatten-$p$ quasi-norms. This connection enables us to minimize the Schatten-$p$ quasi-norm in LRTC implicitly on big tensors and provides an arbitrarily sharper rank proxy for low-rank tensor recovery compared to the nuclear norm.
@@
The Tucker decomposition of a tensor $\bm{\mathcal{T}}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ can be formulated as
$\bm{\mathcal{T}}\approx\bm{\mathcal{G}}\times_{1}\mathbf{U}^{(1)}\times_{2}\mathbf{U}^{(2)}\times_{3}\cdots\times_{d}\mathbf{U}^{(d)}$ (33)
@@
where $\bm{\mathcal{G}}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{d}}$ is the core tensor (a compressed version of $\bm{\mathcal{T}}$), $\mathbf{U}^{(j)}\in\mathbb{R}^{n_{j}\times r_{j}}$ is the factor matrix for mode $j$ (analogous to singular vectors in matrix SVD), and $\times_{j}$ denotes the mode-$j$ product (a tensor-matrix multiplication along mode $j$ ). See (Kolda and Bader, 2009) for more details. The rank of $\bm{\mathcal{T}}$ based on Tucker decomposition is a tuple $(r_{1},r_{2},\ldots,r_{d})$, which is the size of the minimum core tensor $\bm{\mathcal{G}}$ ensuring equality in (33). Examples of Tucker decomposition-based LRTC methods are (Xu et al., 2013; Xu and Yin, 2013; Kressner et al., 2014; Kasai and Mishra, 2016; Xie et al., 2018; Kong et al., 2018).
@@
For convenience, Table 4 summarizes tensor completion approaches based on different decomposition models. The table shows that only a few deep-learning-based methods exist. For instance, Fan (2022a) proposed a multi-mode deep tensor factorization method for tensor completion and provided a generalization error analysis. Their idea is to perform further factorization on the Tucker factor matrices using neural networks. Separately, Ahn et al. (2024) proposed NeAT, which applies neural networks to each latent component in an additive fashion. This approach not only captures diverse patterns and complex structures in sparse tensors but also offers direct and intuitive interpretation by remaining close to the multi-linear tensor model.
@@
The aforementioned models are for regular tensors, where every mode has a fixed size and all slices are aligned. However, there exist irregular tensors in real applications. For example, in a patient-time-feature tensor, each patient may have a different number of visits. In recent years, a few researchers have proposed irregular tensor decomposition and completion methods (Ren et al., 2020; Jang et al., 2022; Kwon et al., 2024; Xie et al., 2025; Han et al., 2025).
@@
(Acar et al., 2011; Jain and Oh, 2014; Zhao et al., 2015)
(Cai et al., 2019; Liu and Moitra, 2020; Fan et al., 2023)
@@
(Gandy et al., 2011; Balazevic et al., 2019; Huang et al., 2015; Xie et al., 2018)
(Chen et al., 2019; Pan et al., 2020; Tong et al., 2022; Wang et al., 2023a)
@@
(Bengua et al., 2017; Ding et al., 2019; Wang et al., 2017; Yuan et al., 2019)
(Huang et al., 2020a; Qiu et al., 2022; Long et al., 2021)
@@
(Zhang and Aeron, 2016; Lu et al., 2018; Zhang et al., 2020a; Jiang et al., 2020)
(He et al., 2022; Wang et al., 2023b; Wu and Fan, 2024)
@@
(Liu et al., 2019; Wu et al., 2019b; Chen and Li, 2020; Tillinghast et al., 2020)
(Fan, 2022a; Ahn et al., 2024; Li et al., 2025b)

### v2

@@ 5.6.4 Graph Neural Network-Based Methods
Representative architectures of GNNs include graph convolutional network (GCN) (kipf2017semi), graph attention network GAT) (velivckovic2017graph), graph isomorphism network (GIN) (xu2019powerful), and graph transformer (GT) (yun2019graph). The hidden layers of a GCN can be formulated as
@@
Several studies have explored using Graph Neural Networks (GNNs) for missing data imputation (zhang2019inductive; you2020handling; spinelli2020missing; telyatnikov2023egg). A key challenge, however, is that GNNs typically require a pre-defined similarity graph across the entire dataset, which is difficult to construct from incomplete data. zhang2019inductive proposed a GNN-based matrix completion method for collaborative filtering. you2020handling proposed GRAPE, a graph-based framework for feature imputation as well as label prediction. GRAPE tackles the missing data problem using a graph representation, where the observations and features are viewed as two types of nodes in a bipartite graph, and the observed feature values as edges. Specifically, in GRAPE, for the incomplete data matrix $\tilde{\mathbf{X}}$ (with mask matrix $\mathbf{M}$) containing $n$ observations and $d$ attributes, a bipartite graph $G=(V_{n}\cup V_{d},E)$ is constructed, where $V_{n}=\{u_{1},u_{2},\ldots,u_{n}\}$ denotes the set of nodes corresponding to the $n$ observations, $V_{d}=\{v_{1},v_{2},\ldots,v_{d}\}$ denotes the set of nodes corresponding to the $d$ attributes, and $E=\left\{\left(u_{i},v_{j},e_{u_{i},v_{j}}\right):u_{i}\in V_{n},v_{j}\in V_{d},e_{u_{i},v_{j}}=\tilde{x}_{ij},M_{ij}=1\right\}$ denotes the edge set. The prediction for the missing values in $\tilde{\mathbf{X}}$ is given by
$\hat{x}_{ij}=f_{\theta}(u_{i},v_{j};G,\mathbf{H}_{n},\mathbf{H}_{d}),\quad\forall(i,j)\text{ with }M_{ij}=0$ (35)
where $f_{\theta}$ is a GNN such as GraphSAGE (hamilton2017inductive). $\mathbf{H}_{n}\in\{1\}^{n\times d}$ is the input embedding matrix for the $n$ sample nodes and $\mathbf{H}_{d}=\mathbf{I}_{d}$ is the input embedding matrix for the $d$ attribute nodes; they are updated in the subsequent layers via neighbor aggregation over $G$. The parameters $\theta$ of the GNN are then optimized using a loss function such as $\sum_{(i,j):M_{ij}=1}(\tilde{x}_{ij}-\hat{x}_{ij})^{2}$.
More recently, telyatnikov2023egg proposed EGG-GAE, a GCN-based method for tabular data that adaptively constructs the adjacency matrix from a neural network’s representations of the incomplete data. In a different domain, li2024stmcdi combined a GNN with a diffusion model to impute missing values in spatial transcriptomics data. More methods of GNN-based imputation can be found in (huang2022graph; wu2023overview; eskandari2024gn2di; zhang2025enhancing).
@@ 5.7 Large Language Model for Missing Data Imputation
Large Language Models (LLMs) (achiam2023gpt; naveed2025comprehensive) are billion-parameter neural networks that learn universal linguistic patterns by next-token prediction on web-scale text. Their recent ascendancy has redefined state-of-the-art across conversational AI, code synthesis, and scientific discovery. It is also promising to exploit LLMs in missing data imputation tasks. Indeed, the pretrained language models are already effective enough in predicting the missing words in pure text owing to the training strategy of next-token prediction, of which a typical application is code completion (wei2023copiloting; izadi2024language; husein2025large). For non-text data such as tabular data, a rudimentary approach is to feed the LLM an incomplete table as raw text and prompts it to hallucinate the blanks; while clever prompting can yield plausible guesses, this tactic barely scratches the surface of LLMs’ expressive power.
[…]
$\mathbf{z}^{(j)}=\text{LM}_{\theta}\left(\text{Textualize}\left(\bar{\mathbf{x}}^{(j)}\right)\right)$ (36)
where $\text{Textualize}(\cdot)$ denotes the operation to convert the input data into text that can be directly handled by the LM. In (9458712), a very simple textualization strategy is used. For example, if $\bar{\mathbf{x}}^{(j)}$ has three attributes including city, hobby, and age, its textualization is [[CLS]; paris; football; 25; [SEP]], where [CLS] and [SEP] are special tokens in BERT (devlin2019bert) or its variants. In (36), $\mathbf{z}^{(j)}$ is the embedding of [CLS] that summarizes the embeddings of the input tokens.
Then IPM fine-tunes the LM and trains a classifier $f^{(j)}_{\phi}$ by
$\mathop{\text{minimize}}_{\theta,\phi}~\frac{1}{|\mathcal{D}|}\sum_{\mathbf{x}\in\mathcal{D}}\text{CE}\left(\text{Softmax}\left(f^{(j)}_{\phi}(\mathbf{z}^{(j)})\right),\mathbf{y}^{(j)}\right)$ (37)
where CE denotes the cross-entropy loss and $\mathbf{y}^{(j)}$ is a one-hot label vector with $y_{k}^{(j)}=1$ if $x_{j}=v_{k}$. Once the models are well trained, the missing attribute $x_{v}$ in a test sample $\mathbf{x}$ is imputed as $v_{k}$ if the $k$-th element of $\text{Softmax}(f^{(j)}_{\phi}(\mathbf{z}^{(j)}))$ is largest. This method is called IPM-Multi. IPM has another variant called IPM-Binary, which models it as a binary candidate selection problem to mitigate over-fitting in low-redundancy datasets. By utilizing fine-tuned models like RoBERTa (liu2019roberta) and neighbor-based negative sampling, IPM captures fine-grained token semantics, outperforming traditional and generative baselines across multiple real-world datasets. One limitation of IPM is that it requires a sufficient number of complete samples to train the classifiers. Another limitation is that it is not effective in inputting numerical data. Note that larger LMs and more sophisticated textualization and embedding summarization methods can be used for (36), which may improve the imputation performance.
@@
More recently, zhang2023large explored the use of LLMs like GPT-4 for data preprocessing tasks such as missing data imputation. The paper introduced a framework that combines prompt engineering techniques—such as zero-shot, few-shot, and batch prompting—with contextualization and feature selection to guide LLMs in imputing missing values. wang2024treb introduces TREB, a BERT-based framework for imputing missing values in tabular data. By transforming numerical rows into tokenized text and fine-tuning RoBERTa with a custom 4-digit tokenizer, TREB captures feature interdependencies without relying on categorical semantics. In (huang2024missing), the authors proposed a missing data imputation framework called PRPMI. PRPMI enhances pre-trained language models by incorporating trainable prompt embeddings and retrieval-augmented knowledge from external data sources. It serializes tabular data into natural language format, fine-tunes the model using both original and retrieved contextual information, and treats imputation as a multiclass classification task. ding2024data proposed to fine-tune a pre-trained GPT-2 model using complete data and prompting it to predict missing values in recommender systems. The method generates semantically meaningful imputations and outperforms a few traditional imputation methods on some benchmarks. hayat2024context presented a method called CRILM for context-aware missing data imputation that leverages pre-trained language models to generate meaningful, contextually relevant descriptors for missing data instead of relying on traditional numerical estimation. By transforming tabular data into natural language and fine-tuning lighter LLMs on these enriched datasets, CRILM effectively handles various missingness mechanisms (MCAR, MAR, and MNAR), and provides promising imputation performance. fang2025spatiotemporal proposed SPLLM, a spatiotemporal pretrained LLM designed for forecasting with missing values. SPLLM integrates a spatiotemporal fusion GCN module to extract spatial and temporal correlations, employs an FFN fine-tuning strategy to adapt pretrained LLMs to incomplete data, and uses a final fusion layer to combine learned and pretrained knowledge.
@@ 5.8 Other Imputation Methods
Some imputation methods fall outside the categories discussed in previous sections. For instance, muzellec2020missing proposed a method based on optimal transport (villani2008optimal). Their core idea is that two randomly drawn batches from the same dataset should share the same distribution; this principle forms a loss function for imputation by matching distributions. Specifically, let $\theta$ be the missing values of the incomplete dataset $\tilde{\mathbf{X}}$, the method solves the following problem
$\mathop{\text{minimize}}_{\theta}~\mathbb{E}_{\tilde{\mathbf{X}}_{A},\tilde{\mathbf{X}}_{B}\sim\mathbf{X}}[\text{Sinkhorn}(\tilde{\mathbf{X}}_{A},\tilde{\mathbf{X}}_{B})]$ (38)
where $\tilde{\mathbf{X}}_{A}$ and $\tilde{\mathbf{X}}_{B}$ are two batches drawn from $\tilde{\mathbf{X}}$ and Sinkhorn is the Sinkhorn divergence (genevay2018learning), an extension of the Wasserstein distance (see (46)). Two variants of (38) were also proposed in (muzellec2020missing).
Inspired by (muzellec2020missing), pmlr-v202-zhao23h proposed imputing missing values by transforming two sample batches into a latent space using deep invertible functions and then matching their distributions. Traditional missing data imputation methods are simple and flexible but require precise model specification, while deep generative approaches are efficient but harder to optimize and rely on stronger assumptions. Motivated by this, jarrett2022hyperimpute introduced HyperImpute, a framework that combines their strengths by adaptively configuring column-wise models. The proposed method is implemented with ready-to-use tools and validated through extensive experiments, demonstrating superior accuracy over benchmarks and reinforcing the effectiveness of iterative imputation.
Interestingly, awan2022reinforcement proposed a reinforcement learning (sutton1998reinforcement; kaelbling1996reinforcement) method for missing data imputation, by learning a policy through an action-reward-based experience. The method was compared with a few baselines such as GAIN and showed competitive performance in the experiments. One limitation of the method is that it performs imputation for each variable separately and hence cannot effectively utilize the correlation between variables.
@@ 6.1 Tensor Completion
Tensors in this paper are referred to as multidimensional arrays (kolda2009tensor) that, as shown by Figure 9, generalize scalars, vectors, and matrices to higher dimensions. We denote $\bm{\mathcal{X}}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ an order-$d$ or $d$-way real tensor. Higher-order ($d\geq 3$) tensors widely exist in science and engineering. Examples of 3rd-order tensors include RGB images and hyperspectral images. Examples of 4th-order tensors include a batch of RGB images and videos. Examples of 5th-order tensors include 3D MRI over time and subjects.
@@
Tensor completion aims to recover the missing entries of a tensor satisfying some properties. Let $\Omega$ be the set of locations of observed entries of $\tilde{\bm{\mathcal{X}}}$. Tensor completion usually solves the following problem
[…]
where $R$ denotes a regularizer ensuring a certain property on $\hat{\bm{\mathcal{X}}}$. Similar to matrix completion, the most common property used in tensor completion is the low-rankness. Low-rank tensor completion (LRTC), as a higher-order extension of LRMC, has been extensively studied in the past decades. A naive approach to LRTC is to unfold a tensor to a matrix first and then perform LRMC or other missing data imputation methods (liu2012tensor). However, unfolding will destroy the higher-order structure of tensors and lead to lower imputation performance. yuan2016tensor pointed out that the unfolding-based tensor completion algorithms require much more observed entries for exact recovery than direct tensor completion without matricization.
LRTC methods can be organized into different categories according to the types of factorization models. Representative tensor factorization models include CP (CANDECOMP/PARAFAC) decomposition (carroll1970analysis), Tucker decomposition tucker1966some, t-SVD decomposition (kilmer2011factorization), tensor train decomposition (oseledets2011tensor), and tensor ring decomposition (zhao2016tensor). In this paper, we take the CP decomposition of a tensor $\bm{\mathcal{X}}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ as an example:
$\bm{\mathcal{X}}\approx\sum_{i=1}^{r}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\circ\cdots\circ\mathbf{a}_{i}^{(d)}$ (40)
@@
$\textup{rank}(\bm{\mathcal{X}})=\min\left\{r\in\mathbb{N}:\ \bm{\mathcal{X}}=\sum_{i=1}^{r}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\cdots\circ\mathbf{a}_{i}^{(d)}\right\}.$
@@
CP decomposition based LRTC methods (acar2011scalable; jain2014provable; zhao2015bayesian; NEURIPS2020_dab1263d; fan2023euclideannorminduced) often solve
$\mathop{\text{minimize}}_{\{\mathbf{a}_{i}^{(j)}\}}~\dfrac{1}{2}\left\|\mathcal{P}_{\Omega}\left(\tilde{\bm{\mathcal{X}}}-\sum_{i=1}^{r}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\circ\cdots\circ\mathbf{a}_{i}^{(d)}\right)\right\|_{F}^{2}+\lambda\cdot R\left(\{\mathbf{a}_{i}^{(j)}\}\right)$ (41)
where the recovered tensor is $\hat{\bm{\mathcal{X}}}=\sum_{i=1}^{r}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\circ\cdots\circ\mathbf{a}_{i}^{(d)}$. The nuclear norm (friedland2018nuclear) of a tensor $\bm{\mathcal{X}}$, though still NP-hard to compute, is defined as
$\|\bm{\mathcal{X}}\|_{\ast}=\inf\left\{\sum_{i=1}^{r}|s_{i}|:\ \bm{\mathcal{X}}=\sum_{i=1}^{r}s_{i}\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\cdots\circ\mathbf{a}_{i}^{(d)},~\|\mathbf{a}_{i}^{(j)}\|=1,\ r\in\mathbb{N}\right\}.$ (42)
@@
$\mathop{\text{minimize}}_{\{s_{i},\mathbf{a}_{i}^{(j)}\}}~\dfrac{1}{2}\left\|\mathcal{P}_{\Omega}\left(\tilde{\bm{\mathcal{X}}}-\sum_{i=1}^{r}s_{i}\cdot\mathbf{a}_{i}^{(1)}\circ\mathbf{a}_{i}^{(2)}\circ\cdots\circ\mathbf{a}_{i}^{(d)}\right)\right\|_{F}^{2}+\lambda\cdot\|\mathbf{s}\|_{1}$ (43)
@@
where $\|\mathbf{a}_{i}^{(j)}\|=1$, $i=1,\ldots,r$, $j=1,\ldots,d$. Note that the $\ell_{1}$ norm in (43) can be replaced by $\ell_{p}$ quasi-norms, leading to tensor Schatten-$p$ quasi-norm regularized LRTC (fan2023euclideannorminduced). In (bazerque2013rank; pmlr-v51-cheng16; yang2016tensor; lacroix2018canonical; fan2023euclideannorminduced), $s_{i}$ are absorbed into the factor vectors $\mathbf{a}_{i}^{(j)}$, which makes the optimization more efficient. In particular, (fan2023euclideannorminduced) presented a class of tensor rank regularizers based on the Euclidean norms of the CP component vectors of a tensor and showed that these regularizers are monotonic transformations of tensor Schatten-$p$ quasi-norms. This connection enables us to minimize the Schatten-$p$ quasi-norm in LRTC implicitly on big tensors and provides an arbitrarily sharper rank proxy for low-rank tensor recovery compared to the nuclear norm.
@@
The Tucker decomposition of a tensor $\bm{\mathcal{X}}\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{d}}$ can be formulated as
$\bm{\mathcal{X}}\approx\bm{\mathcal{G}}\times_{1}\mathbf{U}^{(1)}\times_{2}\mathbf{U}^{(2)}\times_{3}\cdots\times_{d}\mathbf{U}^{(d)}$ (44)
@@
where $\bm{\mathcal{G}}\in\mathbb{R}^{r_{1}\times r_{2}\times\cdots\times r_{d}}$ is the core tensor (a compressed version of $\bm{\mathcal{X}}$), $\mathbf{U}^{(j)}\in\mathbb{R}^{n_{j}\times r_{j}}$ is the factor matrix for mode $j$ (analogous to singular vectors in matrix SVD), and $\times_{j}$ denotes the mode-$j$ product (a tensor-matrix multiplication along mode $j$ ). See (kolda2009tensor) for more details. The rank of $\bm{\mathcal{X}}$ based on Tucker decomposition is a tuple $(r_{1},r_{2},\ldots,r_{d})$, which is the size of the minimum core tensor $\bm{\mathcal{G}}$ ensuring equality in (44). Examples of Tucker decomposition-based LRTC methods are (xu2013parallel; xu2013block; kressner2014low; kasai2016low; 8000407; kong2018t).
@@
For convenience, Table 4 summarizes tensor completion approaches based on different decomposition models. The table shows that only a few deep-learning-based methods exist. For instance, fan2021multi proposed a multi-mode deep tensor factorization method for tensor completion and provided a generalization error analysis. Their idea is to perform further factorization on the Tucker factor matrices using neural networks. Separately, ahn2024neural proposed NeAT, which applies neural networks to each latent component in an additive fashion. This approach not only captures diverse patterns and complex structures in sparse tensors but also offers direct and intuitive interpretation by remaining close to the multi-linear tensor model.
@@
The aforementioned models are for regular tensors, where every mode has a fixed size and all slices are aligned. However, there exist irregular tensors in real applications. For example, in a patient-time-feature tensor, each patient may have a different number of visits. In recent years, a few researchers have proposed irregular tensor decomposition and completion methods (ren2020robust; jang2022accurate; kwon2024compact; xie2025irtf; han2025irregular).
@@
(acar2011scalable; jain2014provable; zhao2015bayesian)
(cai2019nonconvex; NEURIPS2020_dab1263d; fan2023euclideannorminduced)
@@
(gandy2011tensor; balazevic-etal-2019-tucker; huang2015provable; 8000407)
(chen2019non; pan2020low; tong2022scaling; wang2023implicit)
@@
(bengua2017efficient; ding2019low; wang2017efficient; yuan2019tensor)
(huang2020robust; qiu2022noisy; long2021bayesian)
@@
(zhang2016exact; lu2018exact; zhang2020low; jiang2020framelet)
(he2022tensor; wang2023guaranteed; wu2024smooth)
@@
(liu2019costco; wu2019neural; ijcai2020p339; 9338416)
(fan2021multi; ahn2024neural; 10978100)